OpenStudy (anonymous):

give an example with numbers to show why there is no number property that justifies this question 5/n=n/5 please help that (/) over there is a fraction bar.

OpenStudy (anonymous):

pick any number, say 10 then compute \(\frac{10}{5}=2\) and see that is not the same as \(\frac{5}{10}=\frac{1}{2}\)

OpenStudy (anonymous):

ow okie

OpenStudy (anonymous):

thank u again